### v1

# AI Safety in the Eyes of the Downstream Developer: A First Look at Concerns, Practices, and Challenges

Abstract.
Pre-trained models (PTMs) have become a cornerstone of AI-based software, allowing for rapid integration and development with minimal training overhead. However, their adoption also introduces unique safety challenges, such as data leakage and biased outputs, that demand rigorous handling by downstream developers. While previous research has proposed taxonomies of AI safety concerns and various mitigation strategies, how downstream developers address these issues remains unexplored.
This study investigates downstream developers’ concerns, practices and perceived challenges regarding AI safety issues during AI-based software development. To achieve this, we conducted a mixed-method study, including interviews with 18 participants, a survey of 86 practitioners, and an analysis of 874 AI incidents from the AI Incident Database. Our results reveal that while developers generally demonstrate strong awareness of AI safety concerns, their practices, especially during the preparation and PTM selection phases, are often inadequate. The lack of concrete guidelines and policies leads to significant variability in the comprehensiveness of their safety approaches throughout the development lifecycle, with additional challenges such as poor documentation and knowledge gaps, further impeding effective implementation. Based on our findings, we offer suggestions for PTM developers, AI-based software developers, researchers, and policy makers to enhance the integration of AI safety measures.
AI-based Software, AI Safety, Pre-trained Models, Mixed Methods, Empirical Software Engineering.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06copyright: none
1. Introduction
From autonomous vehicles to powerful chatbots, modern society is adopting AI-based software (Martínez-Fernández et al., 2022) in critical domains, including education (Ciolacu et al., 2018), healthcare (Zhou et al., 2018), and transportation (Menze and Geiger, 2015).
With recent advancements in AI models, the expertise and resources required to train a model from scratch have increased (Goyal, 2017; Wei et al., 2024; Patterson et al., 2021). This has led to a paradigm shift, with developers increasingly adopting pre-trained models (PTMs) from platforms such as Hugging Face (Jiang et al., 2023; Tan et al., 2022). Neglecting the safety of AI-based software has harmed individuals and broader societies (inc, [n. d.]). AI safety (Amodei et al., 2016) — avoiding harms by accidental (Amodei et al., 2016) or deliberate (Hernández-Orallo et al., 2020) mis-use — is an important factor to consider during the software development process.
Many recent endeavours from the research community propose taxonomies of AI safety concerns (Amodei et al., 2016; Gao et al., 2024), evaluation methods (Wang et al., 2023), and mitigation approaches (Yi et al., [n. d.]; Roselli et al., 2019; Ji et al., 2023). Additionally, policy makers are also making efforts to provide regulations on AI safety and responsible AI design and development (eur, [n. d.]; iso, [n. d.]). Despite these efforts, how AI-based software developers perceive AI safety concerns, as well as their practices and challenges during the development phase, is not investigated. Due to the complexity of AI models and the knowledge gap between developers and model designers (Nahar et al., 2022; Kim et al., 2017), developers may approach AI safety differently and face various challenges. This is particularly true for smaller-scale company and open-source developers, where fewer resources are provided and regulations might not be enforced. This gap between AI safety frameworks and real-world practices highlights the need for an investigation into developers’ workflows, challenges, and potential shortcomings to better inform researchers and policymakers in supporting AI safety implementation.
This study aims to investigate the concerns that downstream developers have about AI safety, the approaches they adopt, and the challenges they face throughout the entire AI-based software development lifecycle. To achieve this, we conducted a mixed-method study, beginning with interviews with 18 practitioners experienced in developing AI-based software, each lasting approximately one hour. We then performed a thematic analysis to derive answers to our research questions about concerns, practices, and perceived challenges when handling AI safety issues. To complement the qualitative findings, we designed and conducted a survey with 86 practitioners to triangulate our results, providing quantitative evidence on the importance of AI safety concerns, the frequency of different safety-related actions, and the extent to which developers agree on various AI safety challenges. Finally, we compared developers’ concerns against 874 user-reported AI safety issues from the AI Incident Database, identifying gaps in awareness and practices.
Our results show that downstream developers’ AI safety concerns focus on data, model behaviour, and misuse and exploitation concerns. Despite developers’ generally high awareness of AI safety concerns, their actions in the early stages, particularly preparation and PTM selection, are often inadequate. Additionally, their actions are not systematic and vary significantly in comprehensiveness. The primary challenge stems from insufficient policy and regulation guidance, which is scattered across the entire development lifecycle. Other challenges, including lack of standardised benchmarks and evaluation metrics, and insufficient monitoring and maintenance methodologies, further hinder effective AI safety practices. Based on our findings, we provide concrete suggestions to AI-model developers, developers of AI-based software, researchers, and policy makers. The key contributions of our research are as follows.
  • The first mixed-method study revealing downstream developers’ concerns, practices, and challenges regarding AI safety.
  • An analysis of 874 real-world incidents from the AI Incident Database through the lens of developers’ concerns.
  • Concrete suggestions for different stakeholders to improve AI safety practices in software development.
Significance: Software is used to perform increasingly critical functions in our societies, and it increasingly incorporates AI components to do so. Just as they have done for traditional software, software engineers must reason about the safety of the AI-based systems they develop. Although AI researchers and governments have offered guidance on “AI safety”, to date we lack the perspective of the engineers building the systems. Our study gives a first look at their concerns, practices, and challenges, setting the stage for new knowledge and improved knowledge transfer to increase the safety of the AI-based systems on which we depend.
2. Background and Related Work
Here we discuss software engineering practices for AI-based software (section 2.1), and examine the literature on AI safety (section 2.2).
2.1. Engineering Practices for AI-Based Software
In the “Software 2.0” era (Karpathy, 2017), software engineers develop AI-based software (Martínez-Fernández et al., 2022): systems that integrate both hand-coded algorithms and learned components. The development of these learned components differs from traditional software engineering due to its data-driven nature (Khomh et al., 2018). Nahar et al. (Nahar et al., 2022) identified interdisciplinary challenges in communication, documentation, engineering, and process. Amershi et al. (Amershi et al., 2019) introduced a lifecycle based on experiences from Microsoft teams, highlighting three primary deviations from conventional processes: (1) ongoing discovery and management of ML models, (2) model customisation and reuse, and (3) special handling of AI components (e.g., for monitoring, validation, and continuous improvement). The lifecycle proposed for developing AI-based software includes: model requirements, data collection, cleaning, labelling, feature engineering, model training, evaluation, deployment, and monitoring. These differences demand new engineering practices to ensure robust data pipelines (Rupprecht et al., 2020), model reliability (Liu et al., 2023), and safe deployment (Paleyes et al., 2022).
To accelerate the development of AI-based software, engineers re-use and adapt existing pre-trained models (PTMs) (Davis et al., 2023). One major source of PTMs is the HuggingFace platform: with over 1 million models and datasets (Hugging Face, 2025), HuggingFace is the largest and least restrictive PTM registry (Jiang et al., 2022). As with other forms of software re-use (Krueger, 1992), PTM re-use carries with it the challenges of selecting and evaluating a model (Jiang et al., 2023), with many AI-specific details such as missing model metadata (Jiang et al., 2024b), architectural discrepancies (Montes et al., 2022), and conversion errors (Jajal et al., 2024). Industry leaders also emphasise risks in PTM re-use and the broader AI supply chain. For example, Google extended conventional software supply chain governance to catalogue and manage AI artefacts for security, privacy, and compliance (Chaudhuri et al., 2024). HiddenLayer cautioned that rapid AI adoption may outpace the implementation of adequate safety and security controls (HiddenLayer, 2025).
These discussions underscore some of the difficulties of developing AI-based software, whether directly or via re-use. However, literature does not describe how developers actually address these concerns. In this work, we fill this gap by examining challenges and mitigation approaches adopted by PTM users through qualitative and quantitative analysis.
2.2. AI Safety Concerns and Mitigations
2.2.1. What is AI safety?
Any engineered system should be safe within the parameters of its expected use (Anderson, 1992). According to the IEC 61508 safety standard, safety means that a system “functions correctly or fails in a predictable manner”. In the context of software engineering, safety has been a longstanding goal (Lutz, 2000). Safety can be promoted through controlled engineering processes and audits (Lloyd and Reeve, 2009). Others have proposed safety-aware techniques across the engineering cycle, e.g., during requirements engineering (Martins and Gorschek, 2016), design (Leveson and Harvey, 1983), implementation (Chong et al., 2021), validation (Gladisch et al., 2019), and failure analysis (Anandayuvaraj et al., 2024).
By “AI safety” we mean that an AI-based system is safe following the definition in the previous paragraph. Researchers have described many aspects of safety for AI-based systems, including discrimination (Sheng et al., 2021), misinformation, and over-reliance or unsafe use of software (Weidinger et al., 2021). Some authors also include non-accidental risks stemming from malicious use or adversarial attacks, sometimes referred to as AI security risks (Hernández-Orallo et al., 2020). With the rapid advancement and increasing real-world deployment of AI technologies, safety risks are emerging in the delivered software products. A recent systematic literature review on AI safety underscores the gaps in human value alignment, ethical governance, and safety frameworks for AI system implementation (Salhab et al., 2024). Additionally, production-grade chatbots such as ChatGPT and Gemini have raised safety concerns, including policy compliance issues (Hacker et al., 2023), susceptibility to generating or endorsing violent content (Zhuo et al., 2023), and systemic bias (Fleisig et al., 2024).
2.2.2. Realising AI Safety
Researchers have proposed methods for approaching AI safety concerns during the development of AI-based software, including dataset construction methods (Wang and Levy, 2024), evaluation of AI safety risks (Wang et al., 2023), mitigation strategies (Jung et al., 2024; Chen and Liu, 2023), and monitoring strategies (Naveed et al., 2024). Additionally, governance frameworks have been proposed to address AI safety concerns. For instance, Agarwal et al. (Agarwal and Agarwal, 2024) introduced a seven-layer model to standardise fairness assessment, and Lu et al. (Lu et al., 2024) described responsible AI patterns for governance and engineering. Others have surveyed bias in AI systems across scientific disciplines (McGovern et al., 2024), and proposed governance structures such as independent audits (Falco et al., 2021) and ethics-to-practice frameworks (Shneiderman, 2020; Amodei et al., 2016). Additionally, policies like the EU AI Safety Standards (Pouget and Zuhdi, 2024) or ISO 42001 (iso, [n. d.]) aim to regulate high-risk AI systems at a higher level without concrete action suggestions (Pouget and Zuhdi, 2024).
Despite ongoing efforts from AI researchers and policymakers for AI safety, little attention has been paid to understanding the engineering practices of downstream developers. Gaining insights from their perspective is crucial, as it reveals the state of practice and the challenges in addressing AI safety. This understanding helps identify and bridge gaps in knowledge (and knowledge transfer) between AI research, regulations, and real-world software engineering practices.
3. Research Questions & Methodology
We structured our research into three research questions (RQs).
Figure 1. Study design and methodology.
  • RQ1:
    What are software engineers’ safety concerns in AI-based software development?
  • RQ2:
    How do developers approach AI safety issues during the development of AI-based software?
  • RQ3:
    What challenges do developers perceive when handling AI safety issues as they develop AI-based software?
Study Design
To address these RQs, we conducted a mixed-method study, as illustrated in Figure 1. We began by interviewing 18 participants, as interviews are well-suited for exploring complex experiences and eliciting rich, detailed insights from downstream developers (Jiang et al., 2023). This was followed by a survey with 86 participants to validate and quantify the findings (Aniche et al., 2018), and an analysis of 874 incident cases from the AI Incident Database to triangulate our results against real-world AI safety failures (Anandayuvaraj et al., 2024).
In the interview, we did not explicitly define “AI Safety” for the participants. Given the sufficiently abstract concept, we instead asked the concerns they have during development to avoid subjectivity. We further calibrate if the participants’ answers go far beyond general AI safety scopes (Hernández-Orallo et al., 2020), such as carbon emission or super-intelligence replacing humans (“singularity”). In the survey instrument, we asked more concretely about AI safety concerns derived from the interview results.
Our study was conducted with oversight by our institution’s ethics board. Due to its constraints, we cannot share the raw interview transcripts. However, all study instruments, detailed interview analysis, raw survey responses, and all data about AI incident analysis are available in our artefact (section 8).
3.1. Method 1: Interviews
Instrument Design. We designed a semi-structured interview protocol exploring the three research questions. Following the AI-based software development process outlined by Amershi et al. (Amershi et al., 2019), we structured the interview into sections covering background understanding, model selection, model development and testing, model deployment and maintenance. For each stage, we prepared questions asking participants’ development experience, concerns, practices, and challenges. We conducted three pilot interviews with AI-based software developers and then revised our interview protocol based on the results. The final version of our interview took approximately one hour. Table 1 shows sample questions.
Table 1. Summary of interview and survey instruments. (See complete data in replication package)
Topic (# Questions) Sample Questions Survey Scales
Background (5) What is your role in your team? N/A
Concerns (4)
What concerns regarding the poten-
tial AI safety issues would you
consider at model selection stage?
5-point scale: From not
important at all to
Extremely important
Practices (12)
Identifying the AI safety
concern in the deployment
phase, what would you do
next to approach it?
4-point scale: From I
don’t do this at all
to I prioritise this
as a critical task
Challenges (3)
What challenges do you face
when trying to consider and
handle AI safety concerns
in the development phase?
5-point scale: From
Strongly disagree to
Strongly agree
Recruitment. To access downstream developers actively engaged in AI-based software development, we recruited participants from the Hugging Face ecosystem and our professional network, targeting those with AI-based software development experience. Following the design of a previous study to recruit participants from Hugging Face (Jiang et al., 2024a), we extracted all PRO users and members of organisations with over 50 users, retaining only those labelled as company or non-profit, excluding labels such as community and university. Since Hugging Face profiles sometimes link to personal websites, we contacted users via publicly available emails on their personal websites to comply with site policies (git, [n. d.]). This process yielded 98 PRO users and 783 organisational users, totalling 868 unique users after deduplication.
We first contacted PRO users, followed by organisational users, prioritising those from larger organisations based on the number of members. Recruitment and analysis occurred concurrently. In total, we contacted 667 users on Hugging Face. 17 users scheduled interviews with us, of which 13 attended. This resulted in a response rate of 2.55% similar to other software engineering studies (Lethbridge et al., 2005; Singer et al., 2014). To complement this sample, we invited five industry AI-based software developers from our personal network, totalling 18 participants. Table 2 shows the demographics of our interview participants. Each participant was compensated \$20 USD for their time.
Table 2. Interview participant demographics.
ID DL Exp. SE Exp. # AI software Degree Org. Size Domain
P1 1 15 1-5 Bachelor Open-source NLP
P2 6 6 5-10 PhD Small CV
P3 2 4 1-5 Bachelor Open-source NLP
P4 1 2 1-5 PhD Large NLP
P5 2 7 1-5 PhD Open-source CV
P6 12 25 1-5 PhD Small NLP
P7 4 2 10-25 PhD Large NLP
P8 2 2 25+ Master Large CV
P9 6 6 1-5 PhD Large CV, NLP
P10 7 4 5-10 PhD Large CV
P11 4 3 25+ Master Medium CV, NLP
P12 4 4 25+ Master Large CV, NLP
P13 3 2 5-10 Bachelor Large NLP
P14 1 2 5-10 PhD Small NLP
P15 3 3 1-5 Bachelor Medium NLP
P16 1 2 25+ Master Small NLP
P17 4 4 25+ PhD Small NLP
P18 7 7 10-25 PhD Medium NLP
Figure 2. Illustration of our coding process.
Analysis. Following suggestions in software engineering interview studies (Hove and Anda, 2005), the first two sessions were conducted by the first author and a senior researcher, who is also co-author of this study. They discussed the interview flow, after which the first author conducted the remaining sessions. We transcribed the interview recordings using Microsoft 365’s transcription tool, verified manually, and organised in NVivo for analysis. Then we adopted suggested steps from thematic analysis (Braun and Clarke, 2006; Cruzes and Dyba, 2011) for the qualitative analysis stage of the interview transcripts. Two authors, including the first author and the author who participated in two interview sessions and closely monitored the interview transcripts, first read through two interview transcripts, summarising the content as key points related to AI safety concerns, practices, and challenges that corresponded to the three research questions, as suggested to initiate the open coding (Hoda et al., 2012). They then proceeded to assign base level of codes to the extracted key points. Subsequently, we developed higher level abstractions of subthemes by grouping relevant codes together and then the highest level abstractions of themes. Consensus was reached before the first author continued to apply this process to the rest of the transcripts. During the coding process, we constantly compared the emerging codes within and across interview transcript key points, with regular discussions held to verify the code assignment. This iterative process of code development led to continuous adjustment of the codes. Grouping of the codes into themes was conducted concurrently, with each subtheme and theme discussed and adjusted as well. Finally, we derived 29, 100, and 55 codes for the three RQs, respectively. Three authors independently reviewed the coding process and the synthesised subthemes and themes for each research question. The results received agreement among the entire group. Figure 2 displays a sample of the coding process. To determine whether additional participants are needed, we measured the saturation. Following Guest et al.’s (Guest et al., 2006) recommendation, we measured saturation by tracking the cumulative appearance of subthemes in each interview. Saturation in all RQs was achieved after interview 13.
3.2. Method 2: Survey
Instrument Design. After gathering qualitative results for the RQs, we conducted a survey to triangulate our findings with quantitative insights. We designed our survey to investigate the perceived importance of checking AI safety concerns (RQ1), frequency of adopted actions (RQ2), and the agreement on the challenges (RQ3) at a subtheme granularity. Similar to the design in other studies (Aniche et al., 2018), respondents indicated their agreement and experience with statements derived from the interview analysis subthemes. Given that developers’ practices depend on the concern they are trying to address, for RQ2, we asked about the frequency of adopting certain actions separately for each of the three main themes from RQ1. Due to the large number of practice and AI safety concern combinations, we selected practices mentioned by at least half of the participants to minimise fatigue and ensure response quality. Detailed scales are shown in Table 1. We allowed users to respond “I do not know” or “Does not apply” for all questions. The survey included 43 statements and took approximately 15 minutes to complete.
Recruitment. We released our survey on the Prolific platform. To filter participants without prior experience in developing AI-based software, we applied a pre-defined filter from Prolific, requiring participants to be employed in the Information Technology (IT) sector, as in a previous study (Gao et al., 2023). We excluded participants who reported no experience in deep learning, software engineering, or in developing AI-based software. Following Schmidt et al., we asked participants to describe one of their last three AI-based software projects and their roles (Schmidt et al., 2023). Two authors independently reviewed the authenticity of the project and job role description to ensure a high quality of recruited participant (Schmidt et al., 2023). Of the 128 participants recruited, we screened out 27 based on experience questions, and 15 more based on project descriptions, resulting in 86 valid participants. For demographics, 51% had built 1-5 AI-based software projects using PTMs, 26% built 5-10, 13% built 10-25, and 12% built more than 25. For deep learning experience, 47% had 1-2 years, 38% had 3-5 years, 9% had 6-10 years, and 6% had more than 10 years. For software engineering experience, 40% had 1-2 years, 21% had 3-5 years, 21% had 6-10 years, and 20% had more than 10 years.
Analysis. Similar to previous studies (Gao et al., 2023), we converted the responses to a scale of 1 to 4 and applied the Wilcoxon signed-rank test (Rey and Neuhäuser, 2011) with Cohen’s d effect size (Becker, 2000) for pairs of practices under different AI safety concerns to identify whether developers adopt strategies differently when addressing different AI safety concerns. This test is well-suited to our data, given the ordinal nature of the scores and the dependency between paired scores, as they were provided by the same group of participants. In cases of missing data, we included only pairs where both responses were available.
3.3. Method 3: AI Incident Analysis
While interviews and surveys capture downstream developers’ perspectives on AI safety, they do not reveal whether these concerns align with real-world failures. To address this gap, we analysed incidents from the AI Incident Database. By triangulating these perspectives, we identify potential mismatches between developer focus and real-world safety outcomes.
Data Source. We downloaded all 874 incidents from the AI Incident Database, which indexes a collective history of harms or near-harms caused by the deployment of artificial intelligence systems in the real world (inc, [n. d.]). Managed by a board of responsible AI experts, the database compiles public submissions and is recognised as a valuable resource for identifying AI system failures (Anandayuvaraj et al., 2024). Each entry in the database includes a description of the incident.
Analysis. To map the incidents to concerns identified in RQ1, two authors independently annotated each incident based on a specific subtheme failure in AI safety concerns derived from RQ1, assigning one or more violations of AI safety concerns. After two rounds of annotation and discussion, with each round containing 50 incidents, the authors reached a Cohen’s Kappa of 0.71, indicating a good level of agreement (McHugh, 2012). The authors continued to annotate 100 incidents per round independently until the dataset was completed, resolving disagreements as they arose.
4. Results
Here we cover our results, organised by RQ. Table 3 summarises.
Table 3. Summary of our results: sub-themes and frequency by interview participants, survey respondent, and AI incident. The bolded phrases of the sub-themes are used as labels in the figures for survey results (Figures 3, 4 and 5).
RQ Theme Subtheme # Participants Survey Score # AI Incidents
Concerns Data Concerns Safety risks in fine-tuning dataset 7 4.56 66
Safety risks in pre-trained dataset 5 4.52 4
Safety risks in data management 4 4.35 27
Model Behaviour Concerns Data leakage and privacy breaches from the model 8 4.55 18
Model output containing harmful, biased or inappropriate content 13 4.24 283
Reliability and robustness of the model 8 4.27 440
Misuse and Exploitation Concerns Malicious use of the software 5 4.35 242
Irresponsible and unintended use of the software 6 4.21 270
Practices Deeper understanding and assessment for safety concerns Understand dataset content 3 —— ——
Understand model information 5 —— ——
Understand sensitivities for the software 3 —— ——
Regular monitoring and assessment for safety concerns Monitor model input and output 7 —— ——
Collect users feedback regarding AI safety concerns 9 2.86 ——
Develop evaluation dataset for AI model safety 10 3.01 ——
Test and evaluate AI model safety 14 3.07 ——
Implementing technical safety guards Follow security practices in development of AI systems 9 3.24 ——
Ensure dataset quality 9 3.25 ——
Improve safety assurance by iteratively refining models and systems 11 3.04 ——
Filter model input and output 7 —— ——
Maintaining documentation and expert consultation Maintain clear documentation 10 2.93 ——
Cross-functional collaboration for AI safety assurance 9 3.09 ——
Relying on external parties and resources Rely on companies to follow AI safety practice 8 —— ——
Rely on external resources for model quality and content integrity 4 —— ——
Challenges Infrastructure gaps Evaluation and benchmark gap 6 4.09 ——
Monitoring system gap 2 4.20 ——
Technical and model-related difficulties Model interpretability and transparency issue 4 4.17 ——
Performance safety tradeoffs 3 3.72 ——
Model monitoring and maintenance issue 4 4.08 ——
Lack of mature process and methodology Safety policy interpretation gap 6 4.01 ——
Unreliable methodology 7 4.13 ——
Resource constraints Lack of financial and time resource 6 4.00 ——
Lack of human resources 5 4.26 ——
Barriers to technical understanding and information access Technology knowledge gap 3 4.21 ——
Poor documentation 4 4.26 ——
4.1. RQ1: AI Safety Concerns
Developers’ AI safety concerns include three themes, namely (1) data concerns, (2) model behaviour concerns, and (3) miseuse and exploitation concerns. Figure 3 illustrates the RQ1 survey results.
Figure 3. Survey results on importance of AI safety concerns.
Theme 1: Data Concerns relates to data content and management in AI software development. The most discussed subtheme, safety risks in fine-tuning datasets, raised by seven participants, highlights concerns about biased, inappropriate, or personally sensitive content in fine-tuning datasets, which could impact model behaviour (Gong et al., 2023). P11 reflected the privacy aspect by stating: “If I’m training any sort of image generation models, …, I can’t just use someone’s photo…”. In our survey, 97% respondents rated fine-tuning dataset content as “very important” or “extremely important”. Safety risks in pre-trained datasets impacts the PTM behaviour. Similar to fine-tuning datasets concerns, practitioners perceive inappropriate, biased, and privacy-sensitive content within pre-trained datasets as AI safety issues. However, while fine-tuning datasets can be later modified, pre-trained dataset content serves as a proxy for understanding the unintended risks of the PTM. Some participants also noted detecting backdoor data. In our survey, 96% rated checking pre-trained dataset content as “very important” or “extremely important”. Safety risks in data management, concerns risks in storing, accessing, and transferring AI artefacts. For example, P7 noted “whether the storage method is safe, and also this includes storage for the data and the model”. In our survey, 84% respondents consider it “very important” or “extremely important” to establish safe data management process.
Theme 2: Model Behaviour Risks refers to safety risks related to model output. Model output containing harmful, biased, or inappropriate content is the most mentioned subtheme, highlighted by 13 participants. Biased content relates to gender, race, and other factors; harmful content includes offensive or violent outputs; and inappropriate content involves sensitive or erotic topics. In our survey, 78% considered it as “very important” or “extremely important”. Data leakage and privacy breaches from the model occur when outputs resemble private data from fine-tuning or when third-party APIs collect sensitive information. This is a key concern in certain industries like banking, where P7 noted that customers prioritise privacy when using AI-based software. Reliability and robustness of the model concern performance for AI systems in critical domains and its robustness to varied inputs. In the context of LLMs, participants also associate this with hallucinations (Zhang et al., 2023). This concern is raised particularly when the downstream application is in critical domains, such as healthcare, or the AI-component works as an essential role in the software pipelines. Data leakage and privacy breaches, along with model reliability and robustness, received strong support in the survey, with 87% respondents rating them as “very important” or “extremely important”.
Theme 3: Misuse and Exploitation concerns relates to how downstream users engage with AI-based software. Six participants highlighted concerns about irresponsible and unintended use of the software including users probing models with unsafe questions, using the software beyond its intended scope, or over-relying on model outputs. For example, P18 reflected on cases when AI-based software used for medical purposes despite not obtaining certification for such use purposes and not under regulation. Meanwhile, malicious use of the software associates with downstream users deliberatively leveraging the AI-based software to steal training data and model information, or generate harmful content using techniques such as jailbreaking (Wei et al., 2024). In our survey, 81% and 87% respondents rated irresponsible and unintended use, and malicious use of software, as “very important” or “extremely important”, respectively. However, irresponsible and unintended use of software received the lowest score among all concerns at 4.21.
AI incident database analysis result: Our AI incident database analysis highlights real-world AI safety failures from a user-facing perspective. This helps assess whether downstream developers’ awareness and practices are sufficient. Table 3 displays the results. Model reliability and robustness failures are most frequent (50.3% of cases). A representative case involves autonomous driving algorithms running red lights or misidentifying objects, leading to severe consequences. 32.4% of the incidents involve inappropriate, harmful, or biased outputs. Misuse and exploitation failures are also frequent. Irresponsible or unintended use cases or over-reliance on the software accounts for 30.9% of the incidents, making it the third most frequent type of incident. A typical case for this type of failure involves over-reliance on the provided output even without human intervention, such as automating employee layoffs without human oversight. Malicious use makes up 27.7%. Since the AI safety incident database is user-facing, data-related failures are less reported: incidents related to fine-tuning dataset content, pre-trained dataset content, data management, and data leakage and privacy breaches, account for 7.6%, 0.5%, 3.1% and 2% of cases, respectively.
To interpret these data, we compare them to the interview and survey results. “Model reliability and robustness” and “Output containing harmful, biased or inappropriate content” are the most commonly discussed subthemes in interviews. Their reported failure cases ranked second and first, respectively, suggesting that despite high awareness of AI safety, issues in mitigation approaches may exist. In contrast, subthemes under misuse and exploitation concerns were less mentioned in interviews, with irresponsible and unintended use of the software receiving the lowest importance score. However, these failure types still account for a significant portion of incidents, indicating relative short for downstream developers’ attention.
Highlights: Developers’ AI safety concerns focus on data, model behaviour, and misuse/exploitation concerns. Meanwhile, the AI incident database reveals model reliability and robustness issues as the most frequent failures, accounting for over half of cases.
4.2. RQ2: Practices when approaching AI safety
The practices we observed span five themes: (1) deeper understanding and assessment for safety concerns; (2) regular monitoring and assessment for safety concerns; (3) implementing technical safety guards; (4) relying on external parties and resources; and (5) maintaining documentation and expert consultation. Figure 4 quantifies the frequency of practitioners’ strategy choices when approaching three identified AI safety concerns (C1: data concerns, C2: model behaviour concerns, C3: misuse and exploitation concerns).
Figure 4. Survey results on frequency of AI safety practices. (* indicates statistical significance vs. other concerns)
Theme 1: Deeper Understanding and Assessment of Safety Concerns focuses on the preparation phase of developing AI-based software, where developers familiarise themselves with potential risks in AI-related artefacts, development processes, and use cases. Understanding Model Information involves reviewing PTM information related to ethical and robustness aspects from online sources, such as research papers or benchmarks. Additionally, P8 suggested manually inspecting PTM parameter storage and loading scripts to detect malicious code, and also endorsed using the SafeTensors format (hug, [n. d.]) PTMs over other binary formats. For understand dataset content, developers examine both the pre-trained dataset content regarding sensitive and inappropriate content, as a proxy for the PTM behaviour, and fine-tuning dataset to assess data collection quality and potential risks in subsequent AI-based software processes. These manual checks involve reviewing subsets of dataset content or descriptive dataset information from sources such as Hugging Face data cards. Additionally, understand sensitivities for the software involves identifying the target users of the AI-based software and specific sensitive information for them.
Theme 2: Implementing technical safety guards involves adopting methods to mitigate AI safety issues during development. Improve safety assurance by iteratively refining models and systems requires developers to adjust models, evaluate safety factors, and iterate until satisfactory results are achieved. The fine-tuning process may be examined, as P13 mentioned: “first we tried to see if we did something wrong in the fine-tuning process”. This also includes examining fine-tuning processes, switching models or checkpoints, and retraining on improved datasets. For example, P17 fine-tuned with a more diverse accent dataset to enhance speech recognition performance for rural accents. For LLMs, practitioners also refine prompts to prevent undesired content generation. P13 reflected: “in terms of politeness that I mentioned, sometimes we try to change the prompt we are using for the model to prevent it from talking things we don’t want to do”. Our survey revealed significant differences in responses depending on the AI safety context. Developers performed iterative model and system improvement significantly more frequently when addressing model behaviour risks (3.24) compared to data (3.00) and misuse and exploitation (2.98) concerns, with p-values of 0.017 and 0.004, and effect sizes of 0.32 and 0.33, respectively. These results suggest that different AI safety concerns have a medium effect on developers’ frequency of adopting iterative model and system improvement.
The previous efforts relies on the actions to ensure dataset quality. One approach is to incorporate higher-quality data into fine-tuning dataset for the prioritised AI safety concerns. For example, since P17’s speech recognition product targets children in multiple rural areas, they partnered with local organisations to gather diverse accents, enriching their fine-tuning data for model improvement. Filtration methods are also used to remove biased, violent, or personally sensitive content in fine-tuning datasets. For example, when developing AI products for banks, P7 masked customers’ sensitive information before feeding the data into the model, as data privacy is a priority in the financial industry (Akash et al., 2024). These filtration methods involve manual annotation by human reviewers or automated filtering. Also, sampling and data engineering techniques are employed to enhance datasets to provide better inputs aligned with the prioritised AI safety concerns. Our survey shows this practice had the highest average score of 3.25, aligning with the high importance placed on data concerns.
Filtering model input and output acts as an external safeguard against AI safety violations, using both manual rules and automated filters. For example, P15 applied filters to detect inappropriate content, while P12 used regular expressions to remove privacy-sensitive information from model-generated text. For LLMs, P11 highlighted using Llama Guard (Inan et al., 2023) as a post-processing step for model output. Following security practices in the development of AI systems relates to particular processes needed in the development. Several participants highlighted local solutions, such as on-premise data storage and training pipelines to prevent data leakage. For example, P12 addressed data storage concerns by stating: “The data leakage can happen and to resolve it, … on premise the installation of databases”. Other security measures include encrypting data and models, maintaining backups, and selectively publishing checkpoint versions. Based on our survey data, this subtheme is adopted frequently, receiving an average score of 3.24.
Theme 3: Maintaining documentation and expert consultation contains several procedural efforts. Maintain clear documentation refers to the process of providing documentation for both internal review and downstream users to ensure transparency. P9 mentioned that all processes, including decisions on PTM selection, model robustness, and efforts to mitigate risks, need to be clearly documented for company’s audit. Participants also made efforts in recording model evaluations for various AI safety concerns, providing guidelines for responsible interaction with the software, and acknowledging potential AI safety limitations in the documentation. Additionally, documentation on disclaimers was mentioned, with P12 highlighting its use as a way to bypass responsibility for AI safety breaches. Our survey respondents reported varying and slightly lower frequencies of documentation practices across different concerns: data (2.79), model behaviour (3.04), and misuse and exploitation (2.95). Documentation was significantly more frequent for model behaviour concerns than for data concerns (p = 0.017, effect size = 0.30), suggesting a medium impact of concern type on documentation frequency. Overall, documentation ranked among the least frequently adopted AI safety practices.
Meanwhile, cross-functional collaboration for AI safety assurance involves the cooperative efforts to address AI safety concerns, including discussions on risks, collaboration among multiple teams or organisations to collect datasets, test AI models for safety concerns, deploy software securely, and maintenance over time. Moreover, given the expertise gap, AI software developers collaborate with domain experts throughout the development process to analyse risks, build evaluation ground truth datasets, and conduct further evaluations. For example, when handling sensitive customer data, P13 reflected: “I’m not an expert in that field. I just talk with the legal team or the compliance team and if they say no, I just stop … we are working constantly with that team”. This subtheme received an average score of 3.09, indicating practitioners’ broad adoption.
Theme 4: Regular monitoring and assessment for safety concerns involves evaluating AI safety before release and continuous monitoring systems after deployment. A key step is to develop evaluation dataset for AI model safety, with diverse efforts observed, from simply using open-source data to crowdsourcing or even expert consultations. One way to construct evaluation dataset is to reuse existing evaluation dataset on open-source platforms, as P4 mentioned collecting testing documents related to sensitive topics on GitHub. Efforts are also spent refining existing datasets by reformulating certain details. For instance, P6 described adapting culturally American-centric question-answer pairs to reflect the traditions of other culture to evaluate models. When assessing model’s reliability, robustness, or data leakage, practitioners often use ground-truth or production data. For hallucination evaluation, P16 reflected: “the customer support case, …, we will have transcripts of customer support conversations”. Last, P13 also exploit the possibility of using LLM to generate test cases to evaluate model AI safety: “we have used some generative AI models like GPT to maybe help us generate some questions”. Our survey results showed this practice received an average score of 3.01 for its frequency of adoption, indicating it is widely used for AI safety assessments.
After that, developers would test and evaluate AI model safety. Diverse methods were observed among the participants. Some efforts are preliminary, involving only probing the model with small samples and manual inspection,, such as P4:“we randomly select some questions and test the LLM to see if it response with safe answers”. Automatic evaluation using traditional metrics such as Rouge (Lin, 2004) was also mentioned. In particular, LLM-as-judge (Zheng et al., 2023) is used by developers to evaluate model safety factors such as hallucinations, bias, and harmful generation, as reflected by P16. Also, a hybrid approach combines manual and automated effort, as P18 described for privacy evaluation: “then flagging low scoring examples in a development set for … human evaluation”. Our survey indicates that developers conduct model safety testing and evaluation significantly more frequently when addressing model behaviour risks (3.22) compared to data concerns (2.95) and misuse and exploitation concerns (2.94), with p-values of 0.007 and 0.004 and effect sizes of 0.40 and 0.41, respectively. These results suggest that different AI safety concerns have a moderate effect on the frequency with which developers test and evaluate AI model safety.
After deployment, developers made continuous efforts to gauge AI safety concerns in the AI-based software. Regarding monitor model input and output, developers actively track content from both model inputs and outputs to detect any violations of AI safety. For example, P11 noted: “get notification of every prompts that’s been done to my model and I see the output as well”. Moreover, continuous evaluation of the model’s reliability is conducted by periodically measuring key metrics. This allows practitioners to identify model behaviour drift and make necessary adjustments accordingly. Regarding collect users feedback regarding AI safety concerns, practitioners monitor product reviews across various channels, such as app store reviews, and sometimes conduct pre-release user tests. However, this subtheme received the lowest average score of 2.86, with 12.8% of respondents never engaged in this practice. This suggests a relative lack of effort among AI developers in collecting user feedback on AI safety concerns.
Theme 5: Relying on external parties and resources describes developers’ dependence on external components like models, datasets, and libraries for AI safety instead of addressing concerns themselves. In the subtheme Relying on companies to follow AI safety practice, developers often trust the reputation of the companies that deliver AI models. This reliance also extends to company-provided workflows and development processes, such as automated data collection pipelines. Meanwhile, rely on external resources for model quality and content integrity reflects the dependence on popular models, libraries, and datasets.
Highlight: Practitioners’ approaches to AI safety vary in comprehensiveness. Documentation and feedback collection are less frequently adopted, while fewer interview participants consider understanding safety concerns as an initial step.
4.3. RQ3: Perceived Challenges
Five main themes of challenges when handling AI safety concerns are faced by the participants, namely (1) infrastructure gaps, (2) technical and model-related difficulties, (3) lack of mature process, (4) resources constraints, and (5) barriers to technical understanding and information access. Figure 5 displays the survey results.
Figure 5. Survey results on agreement on AI safety challenges.
Theme 1: Infrastructure gaps refers to the absence of artefacts and tools for managing AI safety concerns. Evaluation and benchmark gaps highlights the lack of metrics and benchmarks for evaluating AI safety. P4 noted: “There is no metric to decide if the model is safe enough.”. The lack of benchmark includes both an absence of standardised datasets for specific AI safety concerns and limited datasets marginalised groups such as less common cultures. Similarly, monitoring system gaps reflect the absence of systems to track deployed AI software. For instance, P13 emphasised the need for a dashboard-like monitoring system to proactively detect safety breaches over relying on user feedback. Our survey confirms these challenges, receiving 85% agreement on the challenges of evaluation and benchmark gaps, as well as monitoring system gaps.
Theme 2: Technical and model-related difficulties includes both inherent and external challenges in addressing AI safety issues within model and software systems. Model interpretability and transparency issue hinders model understanding, making safety-oriented fine-tuning and testing difficult. As P12 noted: “AI models are often black boxes, making explainability difficult, which complicates troubleshooting unexpected behaviour.” This was the most acknowledged challenge e.g., with 87% agreement in the survey. Performance safety tradeoffs arises from accuracy loss in safety fine-tuning and efficiency reductions from behavioural guardrails, complicating the balance between performance and safety. This challenge was recognised by 64% of survey participants, though 14% disagreed. The key challenge in model monitoring and maintenance issues is tracking behavioural drift over time and updating systems for AI safety concerns. As P16 noted: “So the questions that are asked, the issues that are faced these change overtime. And so what’s difficult, especially for a customer facing issue is the ability to continue to update the information”.This challenge is worsened by insufficient monitoring metrics and the lack of concrete thresholds for necessary updates. Additionally, adapting updates is difficult due to the challenge of reproducing the observed safety issues. 83% of our survey participants agreed with this challenge.
Theme 3: Lack of mature process refers to insufficient and unreliable methods for addressing AI safety concerns during development. Safety policy interpretation gap highlights the lack of well-defined policies and systematic approaches. For instance, P9 emphasised ambiguity in policy interpretation: “For example, you cannot use untrusted third-party library. The definition is not really clear … so sometimes we need to decide by ourselves.”. P12 noted the complexity in ensuring compliance with legal and ethical standards in multi-regulatory environments. This gap extends to the absence of standardised approaches for managing various AI safety concerns, with P18 adding: “there’s a lack of clarity or even in my opinion, community-agreed best practices on how to do monitoring”. In our survey, 79% respondents agreed that safety policy interpretation gaps pose a challenge. Unreliable methodology refers to the flaws in methodologies themselves, encompassing the limitations of data annotation and crowdsourcing processes (Hsueh et al., 2009). Further concerns include the questionable reliability of user feedback and the inherent difficulty of comprehensively anticipating AI safety issues during development. This challenge was also strongly supported, with 84% of survey participants acknowledging it as a difficulty.
Theme 4: Resources constraints refers to the lack of essential resources. The main subtheme is lack of financial and time resources, as ensuring AI safety requires extra steps, specialised expertise, and more time, creating a significant financial burden. P14 noted: “since we are a small company, we don’t have a big budget”. This challenge was supported by 74% of the survey respondents. With these additional steps for handle AI safety issues also comes lack of human resources. The process of high quality dataset curation requires human resources. Also, most AI-based software developers are not specialised in handling AI safety issues, necessitating the involvement of more experts, as highlighted by P15: “But you don’t have any specialist for the AI safety”. This challenge was agreed by 83% of participants and received the highest score of 4.26, identified as one of the biggest challenges developers face.
Theme 5: Barriers to technical understanding and information access encompasses challenges in accessing and understanding AI safety implementation. The subtheme technology knowledge gaps highlights difficulties arising from unfamiliarity with AI safety terminology and techniques. P2 noted the challenge of technical literacy when discussing AI model privacy metrics: “For our general users, I think most people they do not know the concept of privacy budgets”. This issue also affects cross-team collaboration due to inconsistencies in terminology among stakeholders (Nahar et al., 2022). This challenge is agreed by 86% of the survey participants. Meanwhile, poor documentation refers to incomplete or absent safety evaluation records for PTMs and datasets. P18 highlighted a critical mismatch between available and required safety information: “There’s often a misalignment between what benchmarks are publicly available for the particular task and domains that are actually relevant to the project.” This challenge received the highest average score (4.26), indicating a significant lack of documentation efforts.
Highlights: The lack of human resources and poor documentation are seen as the biggest challenges in addressing AI safety. Meanwhile, gaps in safety policy processing contribute to variations in the comprehensiveness of developers’ practices.
5. Implications
We discuss implications for different stakeholders in this section.
AI-based software developers. Implications are discussed based on the development process (Amershi et al., 2019). (1) Preparation and PTM Selection stage: Based on practices in Table 3, interview participants rarely reflected on deeper understanding and assessment for safety concerns. These actions mainly occur in initial stages of development processes, suggesting insufficient attention from developers in analysing and anticipating AI safety risks from the outset. Meanwhile, PTM attributes and potential risks have been highlighted as essential considerations at the initial stage (Jiang et al., 2023). Therefore, AI-based software developers should dedicate more efforts to assessing risks and intended usage early on, as this proactive approach can reduce downstream challenges and help prevent common misuse and exploitation failures observed in the AI incident database.
(2) Model Development and Testing: Developers frequently adopted practices related to model development and testing, especially for “test and evaluate AI model safety” and “improve safety assurance by iteratively refining models and systems”, which are found even more frequently when dealing with model behaviour concerns in AI safety. However, frequent failures in the AI incident database raise questions about their effectiveness. In fact, we observed varying levels of systematic and comprehensive approaches, with some participants testing AI safety by randomly probing a few examples. Others invested more effort, evaluating real-time data using a combination of automated metrics and human assessment. Recent machine learning research has introduced methods to assess bias (Salewski et al., 2023), toxicity contents (Shaikh et al., 2022), reliability (Wang et al., 2023), with various treatments (Jung et al., 2024; Chen and Liu, 2023). However, knowledge gaps, insufficient documentation, and lack of policies hinder developers from effectively leveraging these methods. To mitigate this knowledge transfer gap, downstream developers should actively seek out new techniques. Given the rapid pace of advancements, aggregator platforms like Reddit or Twitter (Aniche et al., 2018) can serve as valuable resources for staying updated on emerging AI safety solutions. However, developers should also be mindful as not all sources on the Internet are trustworthy, which could cause backfire if adopt blindly.
(3) Model Deployment and Maintenance: Similar to the previous phase, downstream developers mentioned varying practices, though fewer compared to the previous stage were reflected in interviews and the survey (see Table 3 and Figure 4), such as monitor input and output and collect user feedback. This phase is particularly challenging due to limited tool support, unpredictable user interactions, and performance drift. Given the frequent failures for misuse and exploitation in the AI incident database, more frequent and careful practices in this stage are needed. The knowledge transfer gap remains for developers, as AI safety monitoring measures are proposed (Lee et al., 2025), catering for unseen scenarios. Downstream developers still need to more proactively engage in acquiring advancements, by reading papers or using aggregator platforms.
PTM developers. As shown in Table 3, poor documentation is one of the biggest challenges for downstream developers, despite PTM developers’ responsibility to provide clear safety information. Empirical studies on current model cards efforts further highlight issues such as lack of detail and missing information (Gao et al., 2024). To address this, PTM developers should use documentation templates such as model cards (Mitchell et al., 2019) as a good starting point, detailing pre-training data sources, model architecture and training details, clear evaluation results, intended use cases, and mitigation strategies, among others. Moreover, the evolution of open-source projects necessitates continuous documentation updates to ensure clarity and consistency for downstream users (Gao et al., 2025). As PTM versions evolve (e.g., architectural updates, expanded training data), documentation must be promptly updated to maintain clarity for users.
Policy makers. Downstream developers frequently reported a safety policy interpretation gap, citing vague or absent guidelines. For instance, guidelines like “do not use untrusted third-party libraries” remain open to interpretation, while many stages simply lack actionable directives. The absence of concrete guidelines on addressing AI safety issues during each stage leads to varied approaches and confuses downstream developers. Smaller companies and open-source projects face even fewer regulations. For instance, P6, from a large company, must encrypt datasets before transfer, as mandated by company policies, whereas P2, from a small team, relies on a verbal agreement to prevent devices with sensitive data from leaving the office, without formal enforcement. This contrast underscores how organisational structures and policies influence developers’ approaches to data concerns.
Despite efforts to apply AI safety frameworks in regulatory practices (nis, [n. d.]; eur, [n. d.]) and curate question banks for responsible AI risk assessment (Lee et al., 2024), these initiatives emphasise principles over concrete development actions at each stage. An evolving checklist outlining specific actions for addressing various AI safety concerns at different stages of AI software development could help standardise practices and serve as a guideline for developers to mitigate these issues effectively (Braz et al., 2022), particularly in small companies and open-source projects, where safety norms are often lacking. Moreover, drawing an analogy to common criteria evaluation assurance level (EAL, [n. d.]), policy bodies could implement an automated assessment framework to evaluate AI systems across different safety dimensions or based on the aforementioned checklist. AI-based software meeting necessary safety requirements could receive tiered certification indicating its compliance level with specific safety requirements, comparable to ISO 42001 certification (iso, [n. d.]) but lightweight, tiered, and automated. This would help downstream users adopt AI-based software that meet established safety standards. Stronger regulations are also needed to prevent misuse. A frequent issue observed in the AI incident database is that many AI-based software applications are used outside their intended purposes, leading to harmful consequences such as employees being fired solely based on AI decisions or non-medical certified software used for medical purposes.
Researchers. For software engineering researchers, our AI incident database analysis reveals model reliability failures constitute more than half of the reported cases, while failures caused by irresponsible, unintended use or over-reliance of the delivered software is also observed taking 30.9% of the incidents. Furthermore, 15.1% of incidents involve both types of failures, highlighting the widespread premature adoption of unreliable AI-based software under unintended scenarios. Meanwhile, numerous research efforts have proposed solutions for AI safety concerns (Salewski et al., 2023; Chen and Liu, 2023), alongside evolving regulations on AI safety (iso, [n. d.]; eur, [n. d.]). However, with the rapid emergence of new research techniques and policy adjustments, downstream developers often struggle to keep pace, leading to information overload. The frequent occurrence of model behaviour failures and irresponsible usage, coupled with the abundant research and regulations, underscores a significant knowledge transfer gap between AI safety research and practical implementation.
Therefore, research aimed at understanding the challenges that contribute to this knowledge transfer gap between upstream AI researchers and regulators and downstream developers can help bridge this divide. Similar to studies on DevOps education challenges (Fernandes et al., 2022), identifying barriers in AI safety knowledge dissemination could improve its practical adoption in development. Regarding automated tool supports, developing recommender systems or summarisation techniques that process up-to-date policy and research documentation could be a promising direction. For example, TaskNavigator, proposed by Treude et al. (Treude et al., 2014), helps developers navigate software documentation based on different programming tasks. Similarly, such systems could provide targeted policy guidelines and highlight the most relevant evaluation and mitigation solutions based on users’ queries about specific AI concerns at different stages of development. This approach could help standardise AI safety practices in AI-based software development.
6. Threats to Validity
Construct validity: AI safety definitions vary in scope, from direct harm (Amodei et al., 2016) to broader use case concerns (Hernández-Orallo et al., 2020). To capture a comprehensive perspective, we asked developers about their concerns rather than imposing a predefined definition, minimising response bias. For development processes, we referred to several AI-based software development processes (Amershi et al., 2019; Jiang et al., 2023) and synthesised to the context of PTM. The development process is then used to structure our interview protocol, with refinement conducted based on pilot study feedbacks, facilitating systematic investigation for each RQ.
Internal validity: One potential limitation is the relatively small number of interview participants. However, we reached subtheme-level saturation, and no new subtheme emerged with five additional participants. For the survey participants, we have to rely on their self-claimed identity. We mitigate the risk by following the process by Schmidt et al. (Schmidt et al., 2023) to filter out participants whose answer in their developed AI-based software and job roles mismatch to ensure the quality of our survey responses. Another threat lies in the subjectivity of coding and synthesising process. To mitigate this, we conducted a rigorous examination of the transcripts, with three authors reviewing the coding assignments and aggregations. The final themes and subthemes were established through group consensus. Additionally, since interview participants did not necessarily develop the AI systems in the incident database, we did not establish direct links but provided directional discussions.
External validity: First, we recruited participants from Hugging Face in descending order of organisation size, which may over-represent large companies. However, the participants’ current roles involve diverse organisation sizes as indicated by the demographics. Second, we did not filter interview candidates based on the type of models or tasks they develop. As indicated by participants demographics, our results lean more towards NLP-based software developers. Last, given the complexity of AI-based software development, each participant may not reflect in all development stages. To mitigate this, we continued recruitment until saturation was reached across all subthemes.
7. Conclusions
In this paper, we investigated the concerns, practices, and challenges downstream developers face in AI safety in AI-based software development. We conducted a mixed-method study by first interviewing 18 participants to derive qualitative insights for each RQ and triangulated the results by conducting a survey with 86 participants for quantitative perspectives. Additionally, we analysed incidents from an AI incident database to understand real-life AI safety failures to help assess developers’ awareness and approaches. Our findings reveal that despite high awareness of AI safety, developers take limited action during the initial preparation and PTM selection stages. Additionally, knowledge gaps, missing policies, and a lack of tool support contribute to widely varying approaches to AI safety. The frequent occurrence of AI incidents, particularly related to model reliability, harmful and biased content, and misuse and exploitation, raises concerns about the effectiveness of these approaches. We provide concrete implications for different stakeholders, emphasising the need to bridge the knowledge transfer gap between research/policy and downstream developers.
8. Data Availability
Data will be released upon acceptance.